OpenStudy (ksaimouli):

secx cscx=tanx+cotx solve only one side lhs and rhs should be euall plz explain how u did dem=nomenator euall

OpenStudy (cwrw238):

secx cscx = 1/cos x * 1/ sin x = 1 / sinx cosx tanx + cot x = sinx / cosx + cosx / sin x = sin^2 x + cos^2 x / sin x cos x = 1 / sinx cos x so secx cscx=tanx+cotx

OpenStudy (ksaimouli):

no u should get both the equations equall rhs and lhs

OpenStudy (cwrw238):

what i've done is shown that lhs = 1 / sinx cos x and rhs = 1/ sinx cos x therefore lhs = rhs

OpenStudy (ksaimouli):

u should only work one side

OpenStudy (ksaimouli):

sin x / cos x + cos x / sin x ( sin ² x + cos ² x) / (cos x sin x) 1 / (cos x sin x) (1/cos x) (1 / sin x) sec x cosec x RHS sec x cosec x LHS = RHS

OpenStudy (cwrw238):

its like saying if x = a and y = a then of course x = y

OpenStudy (cwrw238):

well both proofs are correct - yours and mine